OpenStudy (anonymous):

find the LCM of each set of polynomials 4w-12 , 2w-6 x^2-y^2 , x^3 +x^2y

OpenStudy (whpalmer4):

To find the least common multiple, whether of numbers or expressions, factor it into its prime factors (those which cannot be further factored). Then combine the highest power of every unique factor seen between all of the items over which you are trying to find the LCM. An example with numbers: LCM of 25 and 45 25: 5*5 45: 3*3*5 LCM is 3*3*5*5 = 225 Check: 25 50 75 100 125 150 175 200 225 45 90 135 180 225 225 is the first number that appears in both lists, so it is the LCM of 25 and 45. With expressions: 2w - 8, 6w-24 2w-8 = 2(w-4) 6w-24 = 2*3(w-4) Therefore the LCM of the two is 2*3*(w-4) = 6w-24 Do the problems and I'll check your work...

OpenStudy (anonymous):

4w-12 = 4(w-3) 2w-6 = 2(w-3)

OpenStudy (anonymous):

is it correct so far???

OpenStudy (whpalmer4):

Yes...

OpenStudy (whpalmer4):

But you didn't factor the 4 all the way...

OpenStudy (whpalmer4):

4=2*2

OpenStudy (anonymous):

4w-12 = 2*2 (w-3) 2w-6 = 2(w-3)

OpenStudy (anonymous):

LCM is 2*2 (w-3) = 4w-12

OpenStudy (whpalmer4):

Correct!

OpenStudy (anonymous):

thanks!!!

OpenStudy (whpalmer4):

What did you get for the other problem?